If you want to delay a pulse of light in a laser experiment, you can send the light through a long coil of fiber optic cable. Light travels somewhat slower in the glass core of a fiber than it does in vacuum. We will approximate the speed of light in the fiber as 2.04 x 10m/s. What length of fiber (in meters) does the light travel in 168 ns?



1
Expert's answer
2021-01-11T03:50:45-0500

Time taken by a certain particle to cover a certain length or distance in particular speed is given by the formula

"Distance = speed \times time"

Here, the particle is light

Speed of light in optical fibre "= 2.04 \times 10^{8} \; m/s"

Amount of time need to be delayed = 168 ns

"= 168 \times 10^{-9} \; s \\

= 1.68 \times 10^{-7} \; s"

The length of optical fibre required = Speed of light in optical fibre x amount of time to be delayed

"= 2.04 \times 10^8 \times 1.68 \times 10^{-7} \; m \\

= 34.27 \;m"

Answer: 34.27 m